OpenStudy (ttop0816):

please help!! Which of the following functions has the same asymptotes as y = cotx? y = cscx y = secx y = tanx y = -tanx

OpenStudy (demandandsupply):

cot x is 1/tan x, therefore the asymptotes are the same when draw in the same graph as y=-tanx @mathmale can you check this and maybe explain why please

OpenStudy (mathmale):

What I'd try first would be to write cot x as cos x cot x = ------- sin x

OpenStudy (mathmale):

Any asymptootes of cot x would be zeros of sin x. Care to experiment with that?

OpenStudy (ttop0816):

then it would be 1/tan!

OpenStudy (mathmale):

Hint: csc x = 1/(sin x)

OpenStudy (ttop0816):

ummm y=secx?

OpenStudy (***[isuru]***):

\[\cot (x) = \frac{ \cos(x) }{ \sin(x) }\] so it has asymptotes wherever sin(x) = 0. What function is the inverse of sin(x)? That will be your answer :)

OpenStudy (ttop0816):

the inverse function would be csc (x) cause its 1/sin(x)?!

OpenStudy (***[isuru]***):

yep!

OpenStudy (ttop0816):

thank you tons!

OpenStudy (***[isuru]***):

u r welcome!

OpenStudy (mathmale):

Note: in "the inverse function would be csc (x) cause its 1/sin(x)" the proper term is "reciprocal," not inverse. The reciprocal of the sine function is the cosecant function. The inverse of the sine function is \[\sin ^{-1}x\], which is very different from the reciprocal. Thought you might want to know this.